OpenStudy (anonymous):

Ronald is 26 years old and starting an IRA (individual retirement account). He is going to invest $200 at the beginning of each month. The account is expected to earn 2.75% interest, compounded monthly. How much money will Ricky have in his IRA when he retires, at age 65?

OpenStudy (anonymous):

pmt = - 200 i = 0.002291667(0.0275/12 as componded mothly) pv =0 n =480( 40*12 as compounded monthly) fv = 174,979.37 amount that ronal will get. Use excel with =fv function and put inputs you`ll get the answers.

OpenStudy (anonymous):

$167,868.30